GachaGirl23:

A rectangle is reduced by a scale factor of One-fourth. A large rectangle has a length of 16 and width of 12. A smaller rectangle has length of 4 and width of 3. Which choices show the ratio of the area of the smaller rectangle to the area of the larger rectangle?

GachaGirl23:

choose three

GachaGirl23:

help please

GachaGirl23:

please

GachaGirl23:

helo?

GachaGirl23:

if yall are not gonna help i am gonaa close the question I really need help with this!

GachaGirl23:

come on guys!!

06natrod52378:

ask @XioGonz

jhonyy9:

so you know area of a rectangle equal length time width smaller rectangle area = 4*3 = ? larger rectangle area = 16*12 = ? ratio = area smaller / area larger @Laylalyssa

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so you know area of a rectangle equal length time width smaller rectangle area = 4*3 = ? larger rectangle area = 16*12 = ? ratio = area smaller / area larger @XioGonZ \(\color{#0cbb34}{\text{End of Quote}}\)